ULTRACOLD MOLECULE FORMATION

\underline{J. T. BAHNS}, W. C. STWALLEY, and H. WANG, Department of Physics, University of Connecticut, Storrs, CT 06269-3046.

Several novel optical and other physical techniques have recently been devised for the control and manipulation of simple diatomic molecules. This talk will review strategies and techniques specifically devised for preparing samples of molecules that are internally as well as translationally untracold.